I tried using the solver for this problem but I still can't get it right. Can someone help me? I have read about legs but I'm not sure which leg is which, so I'll have to descibe the problem as height, width, and diagonal. The problems says:
Use a Pythagorean theorem to find the value of x. The height of the triangle is x-10 and the diagonal is x+10, the width is x. When I used the solver it gave me the answer of 0, which was wrong.
Thanks


x

Found 2 solutions by ankor@dixie-net.com, bigboi:
Answer by ankor@dixie-net.com(22740)

Use a Pythagorean theorem to find the value of x.
The height of the triangle is x-10 and the diagonal is x+10, the width is x.
:
You don't need a solver for something as simple as this:
a^2 + b^2 = c^2 is all you need to know
where a and b are the legs, which is which does not matter
c = the hypotenuse (or diagonal) and is always the longest dimension
:
In your problem
x^2 + (x-10)^2 = (x+10)^2
FOIL
x^2 + (x^2 - 20x + 100) = x^2 + 20x + 100
Combine like terms on the left
x^2 + x^2 - x^2 - 20x - 20x + 100 - 100 = 0
which is
x^2 - 40x = 0
Factor out x
x(x - 40) = 0
Two solutions
x = 0
and
x = 40; the solution we want
:
:
see if this works
a = 40
b = (40-10) = 30
c = (40+10) = 50
40^2 + 30^2 = 50^2
1600 + 900 = 2500; confirms our solution of x=40

Answer by bigboi(1)

x=40 becuase if u replace x for 40 youll get 40^2+30^2= 50^2
so 40^2=1600+30^2 which is 90= square root of 2500
then 1600+90=2500
then square root of 2500 is 50 and c^2=50